Determine the following integral

∫  (1+cos^2 x  / 1+cos 2x ) dx



a.    1 / 2 [ tan x + x] + c


b.   1 / 2[1− 1 / cos x ] + c


c.    1 / 2 + c


d.   1 / 2 tan x + x + c



1
Expert's answer
2022-04-27T17:20:08-0400
"\int \dfrac{1+\cos^2x}{1+\cos2x}dx=\int \dfrac{1+\cos^2x}{2\cos^2x}dx"

"=\dfrac{1}{2}\int \dfrac{1}{\cos^2x}dx+\dfrac{1}{2}\int dx"

"=\dfrac{1}{2}(\tan x+x)+C"

a. "\dfrac{1}{2}(\tan x+x)+C"